OpenStudy (anonymous):

8. An object travels as the description shown below. From t = 0 s to t = 7 s, the object was travelling with a constant velocity of 20 m/s. From t = 7 s to t = 11 s, it underwent a constant acceleration and reached a final velocity of 40 m/s at t = 11 s. The object is travelling to the right for the entire motion. a) What was the acceleration of the object at t = 8 s? b) What was the value of t when the object’s speed was 28 m/s?

OpenStudy (anonymous):

it would be easy to solve these type of problem using velocity-time diagram

OpenStudy (anonymous):

Ya, but he thing is I do not know how to draw the diagram

ganeshie8 (ganeshie8):

it sats acceleration is constant, so from t=7 to t= 11, a is same t = 11-7 = 4 s v = u+ at 40 = 20 + a(4) => a= ?

ganeshie8 (ganeshie8):

*says

OpenStudy (anonymous):